Explain the valuation formula for a constant growth stock.Explain how the formula for a zero growth stock is related to that for a constant growth stock.
1
Expert's answer
2014-12-24T06:36:54-0500
A constant growth stock is a stock whose dividends areexpected to grow at a constant rate in the future. The value of a constant
growth stock can be determined using the following formula
P0 = (D0 ( 1+g))/(r-g)= D1/(r-g)
where
P0 = the stock price at time 0,
D0 = the current dividend,
D1 = the next dividend (i.e., at time 1), g = the growthrate in dividends, and r = the required return on the stock, and g < r.
If there is no growth of dividends g will be equal 0:
g=0
P0 = (D0 ( 1+0))/(r-0)= D/r